OpenStudy (owhemgee):

the energy E stored in an elastic band varies as the square of the extension x. when the elastic is extended by 3cm, the energy stored is 243 joules. what is the energy stored when the extension is 5 cm? What is the extension when the stored energy is 36 joules?

OpenStudy (jack1):

E = C * x^2 Energy = Constant * extension sqrd when x = 3cm, E = 243 so E = C * x^2 243 = C * 3^2 243 = C * 9 243 / 9 = C C = 27

OpenStudy (jack1):

Q1 if E = C * x^2, and x = 5cm, what is E? Q2 if E = C * x^2, and E = 36 joules, what is x?

OpenStudy (jack1):

u following ok @owhemgee ? is this all makin sense dude?

OpenStudy (owhemgee):

yes, ok thank you so much!

OpenStudy (jack1):

all good, u got it from here?

OpenStudy (owhemgee):

yep, thanks man

OpenStudy (jack1):

always welcome dude, slaters